The rate constant for this first‑order reaction is 0.0760s^−1 at 400 degrees Celsius.

A⟶products

After how many seconds will 16.9%of the reactant remain?
1
Expert's answer
2020-02-28T12:13:55-0500

Solution.

"k = \frac{1}{t} \times ln(\frac{С0}{С})"

С0 = 100 %

С = 16.9 %

"t = \frac{ln(\frac{Co}{C})}{k}"

t = 23.39 sec.

Answer:

t = 23.39 sec.